Three point charges are located in the x-y plane.
Answer #1

distance r1=[3^2+4.6^2]^1/2=5.5 cm=0.055 m

the electric field E1=9*10^9*-15.6*10^-9/(0.055)^2=-4.64*10^4 N/c

distance r2=7.8 cm=0.078 m

electric field E2=9*10^9*25.4*10^-9/(0.078)^2=3.8*10^4 N/c

distance r3=0.084 m

electric field E3=9*10^9*85.4*10^-9/(0.084)^2=10.9*10^4 N/c

basing on principle of super position

the net electric field Enet=[10.9+3.8-4.64]*10^4=10.06*10^4 N/c